# arXiv:2603.00659v1[math.FA] 28 Feb 2026

### v1

@@ 1 Introduction
The paper is organized as follows. In Section 2, we recall the definition of Dirichlet form on p.c.f. self-similar set and introduce the Theorem 2.1 and Theorem 2.2. In Section 3, we give the oscillation inequality (OSC) of harmonic function on p.c.f. self-similar set. In Section 4, we prove the Theorem 2.1 and Theorem 2.2. Finally, we give three examples to verify the generalized reverse Hölder inequality (GRH) and Hölder regularity (HR) in Section 5.
@@ 2 Definitions and the Main Result
The analysis on p.c.f. self-similar sets was originally developed by Kigami. For convenience of readers, we will briefly recall the constructions of Dirichlet forms on p.c.f. fractals [26]. Then we give the Theorem 2.1 and Theorem 2.2.
@@ 3 The oscillation inequality of harmonic function
([33, Lemma A.1.]) If $u$ is a harmonic function on $V_{m}$, and $u$ restricted to $V_{0}$ is either 0 or $1$, then the following inequality holds:
@@
By [33, Remark A.2], we know that inequality (3.1) has a clear interpretation in terms of electrical networks. Consider a network in which nodes $x$ and $y$ are connected by a resistor of conductance $H_{x,y}^{w}$. When each boundary point is held at electric potential 0 or $1$, the expression $\mathcal{E}_{m}(u,u)$ equals the total current flowing through the network since a formula $E=IU$ (power equals to current times potential difference). Therefore inequality (3.1) states that the current through any single resistor does not exceed the total network current.
@@
If $u$ is a harmonic function on $V_{m}$, and $u$ restricted to $V_{0}$ is either 0 or $1$, then we have
@@
Similarly, if $u$ is harmonic on $V_{m}$ and $u$ restricted to $V_{0}$ is either 0 or $1$, then
@@
We first provide an introduction to the harmonic function space. Define the space of harmonic functions by $\mathcal{H}\subset C(K)$, which is $N$-dimensional since any harmonic function is uniquely determined by its boundary values. We define the norm of $\mathcal{H}$ by $\|h\|_{\mathcal{H}}^{2}=\mathcal{E}(h,h)+(\sum_{x\in V_{0}}h(x))^{2}$. For every $i=1,\dots,N$, the linear map $M_{i}:\mathcal{H}\to\mathcal{H}$ be defined by $M_{i}h=h\circ F_{i}$. Define $\widetilde{\mathcal{H}}:=\{(h_{1},h_{2},\dots,h_{N-1})\in\mathbb{R}^{N-1}:\sum_{j=1}^{N-1}h_{j}=0\}$.
@@ 5 Example
In ${\mathbb{R}}^{2}$, In $\mathbb{R}^{2}$, let $p_{1}=(0,0)$$p_{2}=(1,0)$, $p_{3}=(1,1)$, $p_{4}=(0,1)$, $p_{5}=\left(\frac{1}{2},\frac{1}{2}\right)$ be the four corners and center of the unit square in the plane. Divide the unit square into a mesh of sub-squares of size $1/9$, and pick $21$ sub-squares as shown in Figure 5. Let $\{a_{i}\}_{i=1}^{21}$ be the center of these sub-squares. Let $\{F_{i}\}_{i=1}^{21}$ be the IFS on $\mathbb{R}^{2}$ with
@@
The Hausdorff dimension of $K$ is $\alpha=\log 21/\log 9$, and the self-similar measure with the natural weight is the normalized $\alpha$-dimensional Hausdorff measure $\mu$ on $K$. The walk dimension of $K$ is $\beta=\frac{\log 21+\log(35/4)}{\log 9}$. By [19, Theorem 6.1], there exist a local regular Dirichlet form on $K$ with the renormalization factor $r=\frac{35}{4}$.
@@
$\left|u(x)-u(y)\right|\leq C\left(\frac{35}{4}\right)^{m}\max_{1\leq k<l\leq 4}\left|u(p_{k})-u(p_{l})\right|.$

### v2

# The Hölder regularity of harmonic function on bounded and unbounded p.c.f self-similar sets

@@ 1 Introduction
The paper is organized as follows. In Section 2, we recall the definition of Dirichlet form on p.c.f. self-similar set and introduce the Theorem 2.1 and Theorem $2.$ In Section 3, we give the oscillation inequality (OSC) of harmonic function on p.c.f. self-similar set. In Section 4, we prove the Theorem 2.1 and Theorem $2.$ Finally, we give three examples to verify the generalized reverse Hölder inequality (GRH) and Hölder regularity (HR) in Section 5.
@@ 2 Definitions and the Main Result
The analysis on p.c.f. self-similar sets was originally developed by Kigami. For convenience of readers, we will briefly recall the constructions of Dirichlet forms on p.c.f. fractals [26]. Then we give the Theorem 2.1 and Theorem $2.$
@@ 3 The oscillation inequality of harmonic function
([33, Lemma A.1.]) If $u$ is a harmonic function on $V_{m}$, and $u$ restricted to $V_{0}$ is either 0 or 11, then the following inequality holds:
@@
By [33, Remark A.2], we know that inequality (3.1) has a clear interpretation in terms of electrical networks. Consider a network in which nodes $x$ and $y$ are connected by a resistor of conductance $H_{x,y}^{w}$. When each boundary point is held at electric potential 0 or 11, the expression $\mathcal{E}_{m}(u,u)$ equals the total current flowing through the network since a formula $E=IU$ (power equals to current times potential difference). Therefore inequality (3.1) states that the current through any single resistor does not exceed the total network current.
@@
If $u$ is a harmonic function on $V_{m}$, and $u$ restricted to $V_{0}$ is either 0 or 11, then we have
@@
Similarly, if $u$ is harmonic on $V_{m}$ and $u$ restricted to $V_{0}$ is either 0 or 11, then
@@
We first provide an introduction to the harmonic function space. Define the space of harmonic functions by $\mathcal{H}\subset C(K)$, which is NN-dimensional since any harmonic function is uniquely determined by its boundary values. We define the norm of $\mathcal{H}$ by $\|h\|_{\mathcal{H}}^{2}=\mathcal{E}(h,h)+(\sum_{x\in V_{0}}h(x))^{2}$. For every $i=1,\dots,N$, the linear map $M_{i}:\mathcal{H}\to\mathcal{H}$ be defined by $M_{i}h=h\circ F_{i}$. Define $\widetilde{\mathcal{H}}:=\{(h_{1},h_{2},\dots,h_{N-1})\in\mathbb{R}^{N-1}:\sum_{j=1}^{N-1}h_{j}=0\}$.
@@ 5 Example
In ${\mathbb{R}}^{2}$, In $\mathbb{R}^{2}$, let $p_{1}=(0,0)$$p_{2}=(1,0)$, $p_{3}=(1,1)$, $p_{4}=(0,1)$, $p_{5}=\left(\frac{1}{2},\frac{1}{2}\right)$ be the four corners and center of the unit square in the plane. Divide the unit square into a mesh of sub-squares of size $1/9$, and pick 2121 sub-squares as shown in Figure 5. Let $\{a_{i}\}_{i=1}^{21}$ be the center of these sub-squares. Let $\{F_{i}\}_{i=1}^{21}$ be the IFS on $\mathbb{R}^{2}$ with
@@
The Hausdorff dimension of $K$ is $\alpha=\log 21/\log 9$, and the self-similar measure with the natural weight is the normalized $\alpha$-dimensional Hausdorff measure $\mu$ on $K$. The walk dimension of $K$ is $\beta=\frac{\log 21+\log(35/4)}{\log 9}$. By [19, Theorem 6.1], there exist a local regular Dirichlet form on $K$ with the renormalization factor $r=\frac{4}{35}$.
@@
$\left|u(x)-u(y)\right|\leq C\left(\frac{4}{35}\right)^{m}\max_{1\leq k<l\leq 4}\left|u(p_{k})-u(p_{l})\right|.$